# Bayesian Model Averaging in Causal Instrumental Variable Models

### v4

@@ 6.3 Many weak instruments
We also consider a setting with many weak instruments similar to the simulation study in Kuersteiner and Okui, (2010). The gIVBMA methods perform best in three out of four scenarios and outperform both IVBMA and state-of-the-art classical estimators. Full details are provided in supplementary Section S.4.3.
@@ 7.1 Geography or institutions?
gIVBMA (BRIC) gIVBMA (hyper-$g/n$) BMA (hyper-$g/n$)
@@
rule 0.85 [0.59, 1.12] 0.87 [0.61, 1.12] 0.79 [0.45, 1.12]
malfal -1.01 [-1.47, -0.54] -1.02 [-1.47, -0.57] -1.05 [-1.67, -0.43]
$\sigma_{12}$ -0.03 [-0.1, 0.03] -0.03 [-0.1, 0.03] - -
$\sigma_{13}$ 0.01 [-0.14, 0.16] 0.01 [-0.12, 0.14] - -
@@
maleco 0.02 1.0 0.01 1.0 0.02 -
lnmort 0.02 1.0 0.01 1.0 0.03 -
frost 0.01 0.03 0.0 0.12 0.02 -
humid 0.02 1.0 0.0 1.0 0.04 -
latitude 0.02 0.13 0.0 0.21 0.01 -
eurfrac 0.03 0.23 0.0 0.61 0.03 -
engfrac 0.01 0.12 0.01 0.15 0.02 -
coast 0.04 1.0 0.0 1.0 0.02 -
trade 0.01 0.21 0.0 0.34 0.02 -
@@
Table 2 shows the results. All outcome PIP are very low, so there is not much evidence for a violation of the exclusion restriction. As expected, maleco and lnmort are relevant instruments. The humidity and coast variables are highly relevant as well. The hyper-$g/n$ prior tends to yield a larger $g$ in the outcome model and a smaller $g$ in the treatment model than the fixed $g$ corresponding to BRIC, leading to lower and higher PIP, respectively. Our treatment effect estimates are in line with those in DiTraglia, (2016). Their point estimate for rule is between $0.81$ and $0.97$, depending on which instrument set they use. For malfal, their point estimates are between 1.16-1.16 and 0.9-0.9. The posterior distributions of both components of the covariance vector, $\sigma_{12}$ and $\sigma_{13}$, concentrate near zero. Thus, the gIVBMA results are similar to naive BMA, though gIVBMA credible intervals are slightly narrower, reflecting the borrowing of strength from the treatment equation.
To assess the predictive performance on this dataset, we compute the LPS using leave-one-out cross-validation. For each observation, the model is trained excluding that observation, and the LPS is then calculated for the excluded data point. We compare the performance of gIVBMA, IVBMA, BMA, and TSLS. Although sisVIVE would be relevant given potential concerns about instrument validity (at least a priori), we exclude it as this approach does not accommodate multiple endogenous variables. Methods designed for many weak instruments were also omitted since their predictive performance was consistently inferior to TSLS in our simulations. Table 3 presents the mean LPS values. The predictive performance of TSLS is competitive with that of gIVBMA, but IVBMA does considerably worse. Naive BMA also predicts very well.
@@
gIVBMA (BRIC) 0.669
gIVBMA (hyper-$g/n$) 0.654
IVBMA (KL) 0.827
BMA (hyper-$g/n$) 0.654
@@ 7.2 Returns to schooling
We revisit this problem using the dataset from Card, (1995), a subset of the National Longitudinal Survey of Young Men, containing all men who were interviewed in 1976 and provided valid wage and education responses. Our outcome is the logarithm of hourly wages with years of schooling as the single endogenous explanatory variable of interest [^3] Card, (1995) also treats experience (and its squared term) as endogenous and uses age and its square as instruments. This conflicts with our distributional assumption as education, experience, and age are perfectly collinear (see supplementary Table S.5), so the joint Gaussian distribution of education and experience would be degenerate conditional on age. Instead, we exclude experience and experience squared and include age and age squared in our set of potential exogenous covariates and instruments. We believe that age sufficiently controls for experience, given that experience is defined as a linear combination of education and age.. Our set of potential exogenous covariates and instruments includes age and age squared, college proximity (distance to a 2 or 4-year college), variables on family background, marital status, race, and regional indicators. The dataset also includes information on parents’ educational attainment. However, due to a substantial proportion of missing values in these variables, we conduct our analysis both including and excluding parental education to assess the robustness of our results. After removing missing values, the data consists of $3,003$ observations without and $2,215$ observations with parental education, and $19$ or $21$ potential instruments and exogenous controls. In both cases, we allow for fully data-driven instrument selection such that all variables can be used as instruments or covariates.
@@
Figure 2: Returns to schooling: Posterior distributions of the treatment effect of education and the covariance ratio $\sigma_{yx}/\sigma_{xx}$ based on the Card, (1995) dataset (a) without parental education ($n=3,003$) and (b) with parental education ($n=2,215$). The algorithm was run for $50,000$ iterations, the first $5,000$ of which were discarded as burn-in. In (b), IVBMA includes a point mass at zero of $6.6\%$ in the posterior distribution of $\tau$.
Figure 2 shows our posterior results [^4] Running the algorithm on this dataset requires only a few hours per specification using a standard 12th Gen Intel i7 CPU. Therefore, applying gIVBMA to datasets with several thousand observations is computationally practical and not a limiting factor.. In line with Card, (1995), the gIVBMA posterior for the treatment effect, $\tau$, indicates higher returns to schooling compared to BMA without endogeneity correction. However, our results vary significantly depending on whether parental education is included. Without parental education, the posterior distribution concentrates slightly above $0.2$, suggesting relatively high returns to schooling. When parental education is included, the posterior mass centers around $0.05$. These latter results indicate lower returns to schooling than those reported by Card, (1995), whose estimates ranged from $0.094$ to $0.194$ across different specifications.
The two datasets lead to different instruments. In the case without parental education, indicators for being Black, living in a Metropolitan area (SMSA), and living in the South consistently appear in the treatment model but rarely in the outcome model, serving as instruments in most analyzed models. When parental education is included, both mother’s and father’s education levels are consistently selected in the treatment model but rarely in the outcome model, making them the primary instruments. The Black and SMSA indicators no longer serve as instruments in this scenario. The college proximity indicators play a less significant role than anticipated. Proximity to a four-year college is occasionally included for the data without parental education, but rarely in the case with parental education, while proximity to a two-year college is rarely included in either case. The IVBMA approach selects college proximity indicators more frequently, using them as instruments in most models without parental education. Complete posterior inclusion probabilities for both cases are presented in Section S.5.1 of the Supplementary Material.
As a robustness check, we re-estimate the model excluding parental education, restricting the sample to observations with complete parental education data ($n=2,215$). The corresponding results are reported in supplementary Section S.5.1. The gIVBMA estimates are broadly consistent with those obtained from the full dataset. These findings suggest that the differences, at least for gIVBMA, are primarily driven by the additional instruments rather than the missing observations, highlighting how posterior inference can be highly sensitive to the choice of the instrument set. We believe the model including parental education provides a better representation of reality. A supporting indicator is the lower estimated variance in the treatment model when parental education is included, indicating a higher degree of instrument relevance. This example serves as a cautionary tale: In a misspecified model (in the sense that important covariates or instruments are missing from the set of available variables), the selection of instruments and the resulting inference on treatment effects can be misleading.
@@
Supplementary Table S.9 shows that the LPS performance of gIVBMA, IVBMA and naive BMA is relatively similar on all three subsets of the data, while TSLS does worse.
@@ 8 Conclusion
We have introduced the gIVBMA method, which performs posterior inference in structural equation models with at least one endogenous regressor, averaging over all possible sets of potential instruments and covariates. We allow for the model to freely choose its instruments based on the data, which provides additional robustness against invalid instruments. Our computational strategy relies on updating the outcome model, the treatment model, and the covariance matrix separately using conditional Bayes factors in the model updates. Exploiting the ULLGM framework (Steel and Zens,, 2025), we can accommodate non-Gaussian outcomes and endogenous variables. We provide simple and unrestrictive necessary and sufficient conditions for model selection consistency of gIVBMA. In both simulation and real data experiments, gIVBMA outperforms the earlier method proposed by Karl and Lenkoski, (2012) and many classical estimators considered in the literature.
@@
All derivations and proofs, along with further details on the prior specification, simulation experiments, and empirical examples, are provided in the Supplementary Material. Replication code is available at https://github.com/gregorsteiner/gIVBMA-Code.
@@
where the instruments $\bm{Z}$ are always excluded from the outcome model. They are, therefore, assumed to be valid a priori. All columns of $\bm{W}$, on the other hand, can still enter either model and can be used either as instruments if they are excluded from the outcome model or as covariates if they enter both models.
@@
The model prior has to be slightly adjusted as the treatment model has more potential variables in this setting. Our choice of the default prior mean model size reflects this by setting $m_{L}=k/2,m_{M}=(k+p)/2$, where $k$ is the number of columns of $W$ and $p$ is the number of columns of $Z$.
@@
S.2.1 Derivation of the conditional outcome distribution
@@
S.2.2 Derivation of the conditional posteriors and marginal likelihoods
S.2.2.1 The Gaussian case
@@
$\quad\cdot\exp\left(-\frac{1}{2}\text{tr}\left(\bm{\Sigma_{xx}^{-1}}(\bm{X}-\bm{V\Lambda})^{\intercal}(\bm{X}-\bm{V\Lambda})\right)\right)\exp\left(-\frac{1}{2}\text{tr}\left(\bm{\Sigma_{xx}^{-1}\Lambda}^{\intercal}(g_{M}^{-1}\bm{V}^{\intercal}\bm{V})\Lambda\right)\right)$
@@
$\quad\cdot\exp\left(-\frac{1}{2\sigma_{y|x}}(\bm{\epsilon}^{\intercal}\bm{\epsilon}-2\bm{\epsilon}^{\intercal}\bm{X}\bm{\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal})-\frac{1}{2}\text{tr}(\bm{\Sigma_{xx}^{-1}B_{\Sigma}}\bm{X}^{\intercal}\bm{X})\right)$
@@
    for $c\in\{l,2l,\dots,(p+k)l\}$. Similarly to the situation for the outcome model, we can easily show that this is always satisfied for the hyper-$g/n$ prior.
@@
S.2.3.3 Extension to the non-Gaussian case
@@
Appendix S.3 Additional Details on the prior specification for $\nu$
Figure S.1 presents different shifted Exponential priors on $\nu$ and their implied priors on the covariance ratio $\sigma_{yx}/\sigma_{xx}$ and the conditional outcome variance $\sigma_{y\mid x}$. The Exponential with scale $1$ strikes a good balance between being uninformative on $\nu$ itself while also keeping the implied prior on the covariance ratio relatively uninformative and is, therefore, our default choice. The latter is almost equivalent to that of fixing $\nu=3$, which is done in Karl and Lenkoski, (2012).
@@
S.4.2 Implementation of competing estimators
@@
  • Ordinary Least Squares (OLS): The OLS estimator is defined as
    $\bm{\hat{\rho}}_{OLS}=(\bm{U}^{\intercal}\bm{U})^{-1}\bm{U}^{\intercal}\bm{y}$
    and its covariance matrix can be estimated as
    $\bm{\hat{\text{var}}}_{OLS}=\hat{\sigma}^{2}(\bm{U}^{\intercal}\bm{U})^{-1},$
    where $\hat{\sigma}^{2}=\left\lVert\bm{y}-\bm{U}\bm{\hat{\rho}}_{OLS}\right\rVert_{2}^{2}/(n-k_{U})$.
@@
  • BMA: Naive BMA refers to using BMA on the outcome model only (and ignoring the treatment equation). We use a $g$-prior specification on the model-specific slope coefficients and uninformative priors on the intercept and the residual variance (see e.g. Fernández et al.,, 2001). Note that this is different from the prior specification of gIVBMA, where the intercept is included in the $g$-prior. BMA is strictly based on the assumption that the outcome is Gaussian. This could be relaxed using the ULLGM approach of Steel and Zens, (2025), but the approximation seems reasonable in our birthweight example, which is the only setting where we consider a non-Gaussian outcome.
@@
  • IVBMA: To implement the IVBMA method by Karl and Lenkoski, (2012), we use the most recent version (1.05, 2014-09-18) of the ivbma R package. This R package is not available on CRAN anymore (withdrawn on 2018-12-07) but can still be downloaded from the archive. Their code does not accommodate the approximate extension to Poisson variables suggested in Lee and Lenkoski, (2022), so when we use IVBMA in the presence of non-Gaussian variables, we use a Gaussian approximation for any non-Gaussian component.

### v5

@@ 6.3 Many weak instruments
We also consider a setting with many weak instruments, similar to the simulation study in Kuersteiner and Okui, (2010), and a variant with a Poisson endogenous variable. In this case, the instruments are assumed to be valid a priori and cannot be included in the outcome model, using the gIVBMA version described in Section S.2. In the Gaussian setting, the gIVBMA variants deliver good coverage and predictions, but are worse than IVBMA in terms of MAE. By contrast, in the Poisson setting, gIVBMA is clearly superior in MAE, and IVBMA has very poor coverage. The horseshoe prior yields lower MAE than gIVBMA in most scenarios, but tends to overcover and predicts poorly. In addition, gIVBMA (like all the other methods with exact zero restrictions) formally deals with identification and has an easier interpretation. MATSLS is competitive for large $n$ but predicts substantially worse than gIVBMA throughout. Full details are provided in supplementary Section S.5.5.
@@ 7.1 Geography or institutions?
gIVBMA (IW) gIVBMA ($\omega_{a}=0.1$) BMA (hyper-g/n)
@@
rule 0.85 [0.6, 1.11] 0.86 [0.61, 1.11] 0.79 [0.45, 1.12]
malfal -1.05 [-1.5, -0.6] -1.03 [-1.46, -0.59] -1.06 [-1.67, -0.43]
$\sigma_{12}$ -0.03 [-0.09, 0.03] -0.03 [-0.09, 0.03] - -
$\sigma_{13}$ 0.02 [-0.11, 0.17] 0.01 [-0.12, 0.13] - -
@@
maleco 0.0 1.0 0.0 0.99 0.02 -
lnmort 0.01 1.0 0.0 1.0 0.02 -
frost 0.01 0.14 0.0 0.16 0.02 -
humid 0.01 0.99 0.0 0.99 0.04 -
latitude 0.0 0.24 0.0 0.21 0.02 -
eurfrac 0.01 0.6 0.0 0.56 0.03 -
engfrac 0.0 0.19 0.01 0.17 0.02 -
coast 0.0 0.99 0.01 1.0 0.03 -
trade 0.0 0.38 0.0 0.42 0.03 -
@@
Table 3 shows the results. All outcome PIPs are very low, while four of the instruments are included in the treatment model with high probability, indicating strong support for overidentification. In particular, maleco and lnmort are valid and relevant instruments, as expected, while the humidity and coast variables are valid and highly relevant as well. The hyper-$g/n$ prior tends to yield a larger $g$ in the outcome model and a smaller $g$ in the treatment model than the fixed $g$ corresponding to BRIC, leading to lower and higher PIP, respectively. Our treatment effect estimates are in line with those in DiTraglia, (2016). Their point estimates for rule and malfal are between $0.81$ and $0.97$ and between 1.16-1.16 and 0.9-0.9, respectively, depending on which instrument set they use. The posterior distributions of both components of the covariance vector, $\sigma_{12}$ and $\sigma_{13}$, concentrate near zero. Thus, the gIVBMA results are similar to naive BMA, though gIVBMA credible intervals are slightly narrower, reflecting the borrowing of strength from the treatment equation.
To assess the predictive performance on this dataset, we compute the LPS using leave-one-out cross-validation. For each observation, the model is trained excluding that observation, and the LPS is then calculated for the excluded data point. We compare the performance of gIVBMA, IVBMA, BMA, and TSLS. Although sisVIVE would be relevant given potential concerns about instrument validity (at least a priori), we exclude it as this approach does not accommodate multiple endogenous variables. Methods designed for many weak instruments were also omitted since their predictive performance was consistently inferior to TSLS in our simulations. Table 4 presents the mean LPS values. The gIVBMA variants perform best, closely followed by naive BMA and TSLS, while IVBMA predicts considerably worse.
@@
gIVBMA (IW) 0.656
gIVBMA ($\omega_{a}=0.1$) 0.654
IVBMA 0.696
BMA (hyper-g/n) 0.658
@@ 7.2 Returns to schooling
We revisit this problem using the dataset from Card, (1995), a subset of the National Longitudinal Survey of Young Men, containing all men who were interviewed in 1976 and provided valid wage and education responses. Our outcome is the logarithm of hourly wages with years of schooling as the single endogenous explanatory variable of interest. Our set of potential exogenous covariates and instruments includes experience, experience squared, college proximity (distance to a 2 or 4-year college), variables on family background, marital status, race, and regional indicators. The dataset also includes information on parents’ educational attainment, which have a substantial proportion of missing values. As in Card, (1995), we impute those missing values using their mean and include indicators for missingness. After imputing missing values, the data consists of $3,003$ observations and $23$ potential instruments and exogenous controls (see supplementary Table S.5). We use fully data-driven instrument selection such that all variables can be used as instruments or covariates.
@@
Figure 2: Returns to schooling: Posterior distributions of the treatment effect of education and the covariance ratio $\sigma_{yx}/\sigma_{xx}$ based on the imputed dataset ($n=3,003$). The algorithm was run for $5,000$ iterations, the first $500$ of which were discarded as burn-in. The vertical line for IVBMA indicates a point mass at zero of $0.962$.
Figure 2 presents our posterior results. [^3] Fitting each specification requires less than an hour on a standard 12th Gen Intel i7 CPU. Thus, computational costs are not a limiting factor for this approach, making gIVBMA practical for datasets with several thousand observations. Consistent with Card, (1995), the gIVBMA posterior for the treatment effect $\tau$ indicates higher returns to schooling than BMA models that lack an endogeneity correction. For comparison, the TSLS estimate when using only proximity to a four-year college as an instrument and all other variables (except proximity to a two-year college) as exogenous controls is $\hat{\tau}=0.142$ ($95\%$ CI: $[0.028,0.256]$). Thus, our posterior results support slightly lower returns to schooling with higher estimation precision, a discrepancy likely driven in part by differences in instrument selection. While our algorithm frequently selects proximity to a four-year college as an instrument, it always includes parental education as well. In contrast, IVBMA yields implausible results by concentrating most of its posterior mass exactly at zero. This might occur because IVBMA fails to select experience in the outcome model and uses it as an instrument instead (see supplementary Table S.6). IVBMA tends to have a higher number of instruments as it selects fewer variables in the outcome model (see supplementary Figure S.6). For all methods, the posterior probability of non-identification (no instruments) is zero.
As a robustness check, we re-estimated the model using only observations with complete parental education data ($n=2,215$; see supplementary Section S.6.1). The results remain qualitatively similar to the main specification. The primary difference is that college proximity is selected as an instrument less frequently, while parental education remains a constant choice. Notably, the IVBMA results appear more plausible in this subsample, as the model shifts toward using parental education as its primary instrument (see supplementary Table S.7).
@@
Supplementary Table S.8 shows that the predictive performance of gIVBMA, IVBMA and naive BMA is relatively similar on both versions of the dataset, while TSLS does worse.
@@ 8 Conclusion
We have introduced the gIVBMA method, which performs posterior inference in structural equation models with at least one endogenous regressor, averaging over all possible sets of potential instruments and covariates. We allow for the model to freely choose its instruments based on the data, which provides additional robustness against invalid instruments. Our computational strategy relies on updating the outcome model, the treatment model, and the covariance matrix separately using conditional Bayes factors in the model updates. Exploiting the ULLGM framework (Steel and Zens,, 2026), we can accommodate non-Gaussian outcomes and endogenous variables. We provide simple and unrestrictive necessary and sufficient conditions for model selection consistency of gIVBMA. In both simulation and real data experiments, gIVBMA outperforms the earlier method proposed by Karl and Lenkoski, (2012) and many classical estimators considered in the literature.
@@
All derivations and proofs, along with further details on the variant with fixed instruments, prior specification, simulation experiments, and empirical examples, are provided in the Supplementary Material. The code and data used are available at https://github.com/gregorsteiner/gIVBMA-Code.
Appendix S.1 Notation
Let $p(\bm{x})$ denote the probability density function (pdf) or probability mass function (pmf) of the random variable $\bm{x}$ and denote by $\mathcal{C}^{q}$ the set of positive definite symmetric (PDS) matrices of dimension $q\times q$. We use $\bm{x}\sim N(\bm{\mu},\bm{\Sigma})$ to indicate that the random vector $\bm{x}\in\Re^{k}$ follows a Gaussian distribution with mean vector $\bm{\mu}\in\Re^{k}$ and covariance matrix $\bm{\Sigma}\in\mathcal{C}^{k}$. We say an $n\times k$ matrix $\bm{R}$ follows a matrix normal distribution, $\bm{R}\sim MN(\bm{M},\bm{S},\bm{T})$ if and only if $vec(\bm{R})\sim N(vec(\bm{M}),\bm{T}\otimes\bm{S}),$ where $\bm{M}\in\Re^{n\times k}$ with $\bm{S}\in\mathcal{C}^{n}$ and $\bm{T}\in\mathcal{C}^{k}$. We use $x\sim IG(a,b)$ to denote that the random variable $x>0$ follows an inverse Gamma distribution with pdf
$p(x\mid a,b)=\frac{b^{a}}{\Gamma(a)}x^{-(a+1)}\exp\left(-\frac{b}{x}\right)$
with shape parameter $a>0$ and scale parameter $b>0$. For a random matrix $\bm{X}\in\mathcal{C}^{k}$, we let $\bm{X}\sim IW(\nu,\bm{\Psi})$ denote that $\bm{X}$ follows an inverse Wishart distribution with pdf
$p(\bm{X}\mid\nu,\bm{\Psi})=\frac{|\bm{\Psi}|^{\nu/2}}{2^{\nu k/2}\pi^{k(k-1)/4}\prod_{j=1}^{k}\Gamma\!\left(\frac{\nu+1-j}{2}\right)}|\bm{X}|^{-(\nu+k+1)/2}\exp\left(-\frac{1}{2}\mathrm{tr}\left(\bm{\Psi}\bm{X}^{-1}\right)\right)$
with degrees-of-freedom parameter $\nu>k-1$ and scale matrix $\bm{\Psi}\in\mathcal{C}^{k}$.
@@
where $\bm{Z}$ is an $n\times p_{1}$ matrix of excluded instruments and $\bm{W}$ is an $n\times p_{2}$ matrix of covariates. The instruments are assumed to be valid a priori. All columns of $\bm{W}$, on the other hand, can still enter either model and can be used either as instruments if they are excluded from the outcome model or as covariates if they enter both models.
@@
The model prior has to be slightly adjusted as the treatment model has more potential variables in this setting. Our choice of the default prior mean model size reflects this by setting $m_{L}=p_{2}/2,m_{M}=(p_{1}+p_{2})/2$.
@@
Appendix S.3 Further details on the prior specification
S.3.1 Prior specification for $\nu$
For the case of the IW prior, Figure S.1 presents different shifted Exponential priors on $\nu$ and their implied priors on the covariance ratio $\sigma_{yx}/\sigma_{xx}$ and the conditional outcome variance $\sigma_{y\mid x}$. The Exponential with scale $1$ strikes a good balance between being uninformative on $\nu$ itself while also keeping the implied prior on the covariance ratio relatively uninformative and is, therefore, our default choice. The priors on the covariance ratio and on $\sigma_{y|x}$ are very close to those implied by fixing $\nu=3$, which is done in Karl and Lenkoski, (2012). For the Cholesky prior, we will choose the same prior on $\nu$.
S.3.2 Prior on the number of valid instruments
An interesting question is what prior is induced on the number of valid and relevant instruments by the prior on the model space described in Subsection 3.3 of the paper. Let $N_{Z}\in\{0,\ldots,p\}$ denote the number of valid and relevant instruments implied by a combination of outcome model and treatment model $(L_{i},M_{j})$, that is, all variables that are included in $M_{j}$, but not in $L_{i}$, denoted as the set $M_{j}\setminus L_{i}$ (in a slight abuse of notation). Then, the distribution of $N_{Z}$ is
$p(N_{Z})$ $=\sum_{i,j:\>\#\lvert M_{j}\setminus L_{i}\rvert=N_{Z}}p(L_{i})p(M_{j})$
$=\frac{1}{p+1}\sum_{p_{i}=0}^{p}\sum_{k=0}^{p_{i}}1_{\{0\leq N_{Z}\leq p-p_{i}\}}\binom{p_{i}}{p_{i}-k}\binom{p-p_{i}}{N_{Z}}\frac{\Gamma(1+p_{i}+N_{Z}-k)\Gamma(1+p-p_{i}-N_{Z}+k)}{\Gamma(p+2)}.$
Figure S.2 presents the prior distribution of $N_{Z}$ induced by our gIVBMA setting for various choices of $p$, and compares it to the prior implied by the IVBMA specification in Karl and Lenkoski, (2012) for $l=1$, which selects a uniform prior over all identified models [^4] This identification condition does not seem to be enforced in the ivbma R package. Instead, it assumes a uniform prior over all models (rather than all identified models). Thus, IVBMA puts some mass on non-identified models, i.e., $N_{Z}=0$, in our simulations and real-data applications.. The main differences are that IVBMA puts zero mass on $N_{Z}=0$ and less prior mass on large values of $N_{Z}$. The IVBMA prior is more uniform on small and moderate $N_{Z}$ for larger $p$, whereas the gIVBMA prior induces a smooth geometric-shaped decline of the prior mass from $N_{Z}=0$ to $N_{Z}=p$ in all cases.
Figure S.2: The implied prior on the number of valid and relevant instruments $N_{Z}$ for our proposed model (gIVBMA) and the one used in Karl and Lenkoski, (2012) (IVBMA; assuming $l=1$) for $p=5,10,15,20$ (from top-left to bottom-right).
@@
S.4.1 Derivation of the conditional outcome distribution
@@
S.4.2 Derivation of the conditional posteriors and marginal likelihoods
S.4.2.1 The Gaussian case
@@
$\quad\times\exp\left(-\frac{1}{2}\text{tr}\left(\bm{\Sigma_{xx}^{-1}}(\bm{X}-\bm{V\Lambda})^{\intercal}(\bm{X}-\bm{V\Lambda})\right)\right)\exp\left(-\frac{1}{2}\text{tr}\left(\bm{\Sigma_{xx}^{-1}\Lambda}^{\intercal}(g_{M}^{-1}\bm{V}^{\intercal}\bm{V})\Lambda\right)\right)$
@@
$\quad\times\exp\left(-\frac{1}{2\sigma_{y|x}}(\bm{\epsilon}^{\intercal}\bm{\epsilon}-2\bm{\epsilon}^{\intercal}\bm{X}\bm{\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal})-\frac{1}{2}\text{tr}(\bm{\Sigma_{xx}^{-1}B_{\Sigma}}\bm{X}^{\intercal}\bm{X})\right)$
@@
    for $c\in\{l,2l,\dots,pl\}$. Similarly to the situation for the outcome model, we can easily show that this is always satisfied for the hyper-$g/n$ prior.
@@
S.4.3.3 Extension to the non-Gaussian case
@@
S.5.2 Implementation of competing estimators
@@
  • BMA: Naive BMA refers to using BMA on the outcome model only (and ignoring the treatment equation). We use a $g$-prior specification on the model-specific slope coefficients and uninformative priors on the intercept and the residual variance (see e.g. Fernández et al.,, 2001). Note that this is different from the prior specification of gIVBMA, where the intercept is included in the $g$-prior.
@@
  • IVBMA: To implement the IVBMA method by Karl and Lenkoski, (2012), we use the most recent version (1.05, 2014-09-18) of the ivbma R package. This R package is not available on CRAN anymore (withdrawn on 2018-12-07) but can still be downloaded from the archive. As no intercept is included by default, we centre the outcome and treatments before using the IVBMA function. Their code does not accommodate the approximate extension to Poisson variables suggested in Lee and Lenkoski, (2022), so when we use IVBMA in the presence of non-Gaussian variables, we use a Gaussian approximation for any non-Gaussian component.
  • BayesHS: We fit a Bayesian IV model with all instruments included and put a Horseshoe prior (Carvalho et al.,, 2009) on both the outcome coefficients $\bm{\beta}$ and the treatment coefficients $\bm{\Delta}$. The horseshoe prior on $\bm{\beta}$ (analogously on $\bm{\Delta}$) corresponds to
    $\beta_{i}\sim N\left(0,\lambda_{i}^{2}\mu^{2}\right),\quad\mu\sim C^{+}(0,1),\quad\lambda_{i}\sim C^{+}(0,1),$
    where $C^{+}$ denotes a half-Cauchy distribution. We refer to the $\lambda_{i}$ as the local and $\mu$ as the global shrinkage parameters and we choose them independently for the outcome and treatment model. We approximate the posterior distribution using variational inference in Turing.jl (Fjelde et al.,, 2025). Our variational family is a Gaussian mean field and we perform $500$ optimisation steps of the distance over weighted gradients algorithm (Khaled et al.,, 2023), which is the default optimiser in Turing.jl. We have also tried sampling from the posterior using a No-U-Turn sampler. However, the convergence of the chains was poor in many cases, and the execution was very slow.